1 pts Question 1 What would be the pH needed to initiate precipitation for a solution that contains 4.88 x 10 M Pb (NO3)2? Ksp of Pb 2 is 1.2 x10 15 (OH)
Answer #1

Kspof Pb(OH)2-1.2x10 [PS-]= 4.88x10-3M 2_ 1.2x10-15 4.88x10M OH 1-0.246x 10-12 [OH-]= 0.496×10-6 pOH =-log(0.496 x 10-6) = 6.